OpenStudy (anonymous):

hi some of my theses questions got qron in multiple choices can you help ____ two points are collinear. A. Any C. No

OpenStudy (anonymous):

answer is any

OpenStudy (anonymous):

any two points are collinear because u can join the two points to form a straight line

OpenStudy (anonymous):

What is a counterexample for the conjecture? Conjecture: The product of two positive numbers is greater than the sum of the two numbers. A. 3 and 5 B. 2 and 2 D. There is no counterexample. The conjecture is true.

OpenStudy (anonymous):

can anyone please help me with some multiple questions i got wrong out of 600 questions

OpenStudy (anonymous):

i think the answer is A

OpenStudy (mertsj):

The conjecture is: multiply two positive numbers. Add the same two numbers. The first answer will always be greater.

OpenStudy (mertsj):

I don't think that is true because 1 times 1 = 1 1+1 = 2. The sum is greater. That is a counter example.

OpenStudy (mertsj):

In your problem they suggested 3 and 5 but that is not a counter example because 3 x 5 = 15 3+5 = 8 15 IS greater so that supports the conjecture.

Parth (parthkohli):

However, 2 and 2 also qualify as counterexamples.

OpenStudy (mertsj):

On the other 2 x 2 = 4 2+2 = 4 so the product is NOT greater and that IS a counterexample.